Jenny and her cousin Jason share the same birthday. Jenny realized that on their next birthday, she would be 2 years less than 3 times Jason's age. If the sum of their ages will be 18 years, how old will Jenny and her cousin be on their next birthday?
Answer by TimothyLamb(4379)

---
x = Jenny
y = Jason
---
x + 1 = 3(y + 1) - 2
x + 1 + y + 1 = 18
---
put the system of linear equations into standard form:
---
x + 1 = 3y + 1
x + y + 2 = 18
---
x - 3y = 0
x + y = 16
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
their ages today:
x = Jenny = 12
y = Jason = 4
---
their ages on their next birthday:
x = Jenny = 13
y = Jason = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php